### v3

# MSC 30C45, 30C50

MSC 30C45, 30C50
WITHDRAWN: A PROOF OF THE KRZYZ CONJECTURE
D. L. Stupin [^1] This is an English translation of a preprint originally published in Russian:
https://preprints.ru/article/1799 (Preprints.ru, 2025).
WITHDRAWN: The proof contains an uncorrectable gap in the proof of theorem 7 on page 11.
A proof of the Krzyz conjecture is presented, based on the application of the variational method, as well as on the use of two classical results and some of their consequences. The mentioned results are the Caratheodory–Toeplitz criterion of continuing a polynomial to a Caratheodory class function, and the Riesz–Fejer theorem about trigonometric polynomials.
This is an English translation of a preprint originally published in Russian: https://preprints.ru/article/1799 (Preprints.ru, 2025).
Keywords: Caratheodory-Toeplitz criterion, Fejer-Riesz Theorem, trigonometric Polynomials, polynomials with positive real part, Caratheodory class, the Krzyz conjecture, the Krzyz hypothesis, the Krzyz problem, bounded nonvanishing functions
Introduction
Let us denote the unit disk by $\Delta\coloneq\{z\,:\,z\in\mathbb{C},\leavevmode\nobreak\ |z|<1\}$, and the unit circumference by $\partial\Delta$. The Taylor coefficients of a function $f$ will be denoted by $\{f\}_{n}$, $n\in\{0\}\cup\mathbb{N}$.
The class $B$ is the set of functions $f$ holomorphic in the unit disk $\Delta$ and satisfying $0<|f(z)|\leqslant 1$, $z\in\Delta$.
In 1968, Jan Krzyz proposed a conjecture [1] that if $f\in B$, then
|$|\{f\}_{n}|\leqslant 2/e,\qquad n\in\mathbb{N},$
with equality attained only for functions of the form $e^{i\theta}F(e^{i\varphi}z^{n},1)$, where
$F(z,t)\coloneq e^{-t\frac{1-z}{1+z}},\qquad\varphi,\theta\in\mathbb{R},\quad t\in[0,+\infty).$ (1)
Fix $n\in\mathbb{N}$. The problem of finding
$m_{n}\coloneq\max\limits_{f\in B}{|\{f\}_{n}|},$
will be called the Krzyz problem for index $n$. A function $f\in B$ will be called extremal in the Krzyz problem for index $n$ if |$|\{f\}_{n}|=m_{n}$.
The problem of finding sharp estimates for |$|\{f\}_{n}|$ for all $n\in\mathbb{N}$ on class $B$ will be called the Krzyz problem. The Krzyz problem for fixed $n$ may also be called simply the Krzyz problem if this doesn’t lead to confusion.
The existence of extremals in the Krzyz problem for fixed index $n$ is obvious, since after adding the function $f(z)\equiv 0$ to class $B$ we get a compact family (in the topology of locally uniform convergence), and the functional assigning to each function from $B$ its $n$-th Taylor coefficient is continuous on $B$.
By $\Omega_{0}$ we denote the class of functions $\omega$ holomorphic in $\Delta$ and satisfying $|\omega(z)|<1$, $z\in\Delta$, $\omega(0)=0$.
Let mappings $G$ and $g$ be holomorphic in $\Delta$. A function $g$ is called subordinate to function $G$ in the unit disk $\Delta$ if it can be represented as $g(z)=G(\omega(z))$, where $\omega\in\Omega_{0}$. In this case, function $G$ will be called a majorant for $g$ in $\Delta$. The theory of subordination is presented in [3].
The class of functions with positive real part in $\Delta$ is denoted by $C$. This class will be called the Caratheodory class. Fix $t>0$. The set of functions $h$ from $C$ normalized by $h(0)=t$ is denoted by $C_{t}$ and will be called the normalized Caratheodory class, or simply the Caratheodory class if no confusion arises. Note that $C\coloneq\bigcup\limits_{t>0}C_{t}$.
Since class $B$ is invariant under rotations in the $w$-plane ($w=f(z)$), we can restrict ourselves to studying functions with $f(0)>0$ without loss of generality. As $0<\{f\}_{0}\leqslant 1$, we can set $\{f\}_{0}=e^{-t}$, where $t\in(0,+\infty)$. The corresponding subclasses are denoted by $B_{t}$. From the theory of subordinate functions [3], each function in class $B_{t}$ can be represented as
$f(z)=e^{-t\,h(z)},\qquad h\in C_{1}.$ (2)
Note that for each $t>0$ this establishes a bijection between $C_{1}$ and $B_{t}$. Clearly, for each fixed $t>0$, the function $F(z,t)$ is a majorant for every function in class $B_{t}$.
Geometrically, it’s obvious that |$|\{f\}_{0}|\leqslant 1$. The sharp estimate for |$|\{f\}_{1}|$ appears in many works since 1934, starting with [2]. The estimate for |$|\{f\}_{2}|$ has been known since 1943 [3]. Krzyz formulated his conjecture based on these estimates. The case $n=3$ was first proved in 1977 by J. Hummel, S. Scheinberg and L. Zalcman [4]. A sharp estimate for |$|\{f\}_{3}|$ on class $B_{t}$ was obtained in [5]. A similar estimate for functions with real coefficients was given in [6]. For $n=4$, W. Szapiel’s proof [7] should be mentioned. The estimate for the fifth coefficient using Szapiel’s method was first given by N. Samaris [8]. In [9], the author obtained |$|\{f\}_{6}|\leqslant 2/e+0.00116077$ using Szapiel’s method, and in [10] - the estimate |$|\{f\}_{6}|\leqslant 2/e$ using numerical methods.
Research on the Krzyz problem can be divided into several directions. One concerns estimating initial coefficients, as mentioned above. Asymptotic estimates, discussed in [11, 12], are another direction. There are also uniform estimates for all $n\in\mathbb{N}$ using Cauchy’s integral formula:  [13] obtained |$|\{f\}_{n}|\leqslant 0.999877\ldots$, while [14, p. 19] improved this to $0.999178\ldots$ Another direction studies properties of extremal functions and searches for functions in $B$ with these properties, as in [4, 15, 16] and the present work.
S.L. Krushkal in the work [17] obtained a proof of the Hummel-Scheinberg-Zalcman conjecture [4, p. 189] for functions from the space $H^{p}$. This proof was obtained by applying a new method based on utilizing deep features of Teichmuller spaces. The Krzyz conjecture is a consequence of the Hummel-Scheinberg-Zalcman conjecture when $p\to\infty$.
Here we present a new elementary proof of the Krzyz conjecture, presented in a self-contained form. All literature references are given for additional reading or as sources.
1. The Caratheodory-Toeplitz criterion
For future reference, we state the following classical result [18, 19]:
Theorem 1 (Caratheodory, Toeplitz).
Let $n\in\mathbb{N}$, $\{h\}_{0}>0$, $\{h\}_{1},\ldots,\{h\}_{n}\in\mathbb{C}$. The polynomial
$Q_{n}(z)\coloneq\{h\}_{0}+\sum\limits_{k=1}^{n}{\{h\}_{k}z^{k}}$ (3)
can be extended to a function
$h(z)\coloneq Q_{n}(z)+o(z^{n})\leavevmode\nobreak\ \in\leavevmode\nobreak\ C$
if and only if the determinants
$M_{k}\coloneq\begin{vmatrix}2\{h\}_{0}&\{h\}_{1}&\cdots&\{h\}_{k-1}&\{h\}_{k}\\ \overline{\{h\}}_{1}&2\{h\}_{0}&\cdots&\{h\}_{k-2}&\{h\}_{k-1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \overline{\{h\}}_{k-1}&\overline{\{h\}}_{k-2}&\cdots&2\{h\}_{0}&\{h\}_{1}\\ \overline{\{h\}}_{k}&\overline{\{h\}}_{k-1}&\cdots&\overline{\{h\}}_{1}&2\{h\}_{0}\\ \end{vmatrix},\qquad k=1,\ldots,n,$ (4)
are either all positive, or positive up to some index $m\leqslant n$ after which they all equal zero. In the latter case, the extension is unique and there exist numbers $\alpha_{k}>0$, $k=1,\ldots,m$, and $0\leqslant\varphi_{1}<\ldots<\varphi_{m}<2\pi$ such that
$h(z)=\sum\limits_{k=1}^{m}\alpha_{k}\frac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}.$ (5)
Fix $n\in\mathbb{N}$ and consider the space $\mathbb{C}^{n+1}$ whose points are $(n+1)$-tuples of complex numbers $h^{(n+1)}\coloneq(\{h\}_{0},\ldots,\{h\}_{n}).$ The set of points $h^{(n+1)}$ in $\mathbb{C}^{n+1}$ such that the numbers $\{h\}_{0},\ldots,\{h\}_{n}$ are the first $n+1$ coefficients of some function $h\in C$ will be denoted by $C^{(n+1)}$ and called the $(n+1)$-th coefficient body of class $C$.
The coefficient problem for class $C$ is formulated as follows: find necessary and sufficient conditions for numbers $\{h\}_{0},\{h\}_{1},\{h\}_{2},\ldots$ such that the series $\{h\}_{0}+\{h\}_{1}z+\{h\}_{2}z^{2}+\ldots$ is the Taylor series of some function in class $C$. The Caratheodory-Toeplitz criterion gives a complete solution to this problem.
2. Some properties of the extremal function
The class $B$ is invariant under rotations in the $z$-plane and rotations in the $w$-plane ($w=f(z)$). That is, if $n\in\mathbb{N}$ and a function $f$ is extremal in the Krzyz problem for this index $n$, then the function $\eta f(\zeta z)$ with $|\eta|=|\zeta|=1$ is also extremal. Here, rotation in the $z$-plane does not affect the coefficient $\{f\}_{0}$. Therefore, without loss of generality, we may assume that if $f$ is extremal, then $\{f\}_{0}>0$ and $\{f\}_{n}>0$. In what follows, when speaking of a function $f$ extremal in the Krzyz problem, we will imply that $\{f\}_{0}>0$, $\{f\}_{n}>0$, and $f\in B_{t}$, where $t=-\ln\{f\}_{0}$. Also, a function extremal in the Krzyz problem for index $n$ will often be called simply extremal if it is clear from the context which specific $n$ is meant.
Fix $n\in\mathbb{N}$. If $f(z)=e^{-h(z)}$ is an extremal function, then, as is known [19, 20], the point $h^{(n+1)}$ belongs to the boundary of the set $C^{(n+1)}$, which is equivalent to $M_{n}=0$ ($M_{n}$ is defined by formula (4)). According to the Caratheodory-Toeplitz criterion, this means that the extension of the polynomial $Q_{n}$, defined by formula (3), is unique. Consequently, any extremal function has the form $f(z)=e^{-h(z)}$, where $h$ is given by formula (5). Thus, the following statement holds:
Corollary 1.
Let $n\in\mathbb{N}$, let $f$ be extremal in the Krzyz problem, and let $t\coloneq-\ln\{f\}_{0}$. Then there exist numbers $\alpha_{k}>0$, $k=1,\ldots,m$, $\sum\limits_{k=1}^{m}\alpha_{k}=t$, and numbers $0\leqslant\varphi_{1}<\ldots<\varphi_{m}<2\pi$, such that $f(z)=e^{-h(z)}$, where the function $h$ is given by formula (5), with $m\leqslant n$.
This result has been well known since the early 20th century (see, e.g., [4, p. 171] and [16, p. 725]).
Note that if $n\in\mathbb{N}$, and $f$ and $g$ are functions holomorphic in $\Delta$, then
$\{f\cdot g\}_{n}=\{f\}_{n}\{g\}_{0}+\{f\}_{n-1}\{g\}_{1}+\ldots+\{f\}_{0}\{g\}_{n}.$ (6)
Lemma 1.
Let $n\in\mathbb{N}$, $f$, $g$ be functions holomorphic in $\Delta$, $\varepsilon\in\mathbb{R}$, and
$v(z)\coloneq f(z)e^{-\varepsilon g(z)},$ (7)
then
$\{v\}_{n}=\{f\}_{n}-\varepsilon\{f\cdot g\}_{n}+o(\varepsilon),\qquad\varepsilon\to 0.$ (8)
Proof. Consider the variation $v$ of $f$ by the function $e^{-\varepsilon g}$, $\varepsilon\in\mathbb{R}$. Letting $\varepsilon$ tend to zero, we have:
$v(z)\coloneq f(z)e^{-\varepsilon g(z)}=f(z)(1-\varepsilon g(z)+o(\varepsilon))=f(z)-\varepsilon f(z)g(z)+o(\varepsilon).$
[…]
The following result is borrowed from [16, p. 726]. For completeness, we present it along with the proof, also taken from that work.
Theorem 2.
Let $n\in\mathbb{N}$ and $f$ be a function extremal in the Krzyz problem, then for any function $g\in C$
$\operatorname{Re}{(\{f\}_{n}\{g\}_{0}+\{f\}_{n-1}\{g\}_{1}+\ldots+\{f\}_{0}\{g\}_{n})}\geqslant 0.$ (9)
In particular, if $h\coloneq-\ln f$, then
$\operatorname{Re}{(\{f\}_{n}\{h\}_{0}+\{f\}_{n-1}\{h\}_{1}+\ldots+\{f\}_{0}\{h\}_{n})}=0.$ (10)
Let
$H(z)\coloneq\{f\}_{n}+2\{f\}_{n-1}z+\ldots+2\{f\}_{0}z^{n},$ (11)
then
$H\in C.$ (12)
Moreover, let $m\leqslant n$, and let the numbers $0\leqslant\varphi_{1}<\ldots<\varphi_{m}<2\pi$ partially define $h\coloneq-\ln f$ (see formula (5)), then
$\operatorname{Re}{H(e^{i\varphi_{k}})}=0,\qquad k=1,\ldots,m.$ (13)
Proof. 1. Let us prove formula (9). Let $g\in C$. The variation $v$ of $f$, given by formula (7), for $\varepsilon>0$ is internal, i.e., $v\in B$. As $\varepsilon\to 0$, Lemma 1 holds, and by assumption $\{f\}_{n}>0$ and $f$ is extremal, hence $\operatorname{Re}{\{v\}_{n}}\leqslant\{f\}_{n}$. From this (see (8)) follows the validity of formula (9).
2. Let us prove formula (10). Now take $g=h$. The variation $v$ of $f$, given by formula (7), for sufficiently small $\varepsilon<0$ is an internal variation, i.e., $v=f\cdot f^{\varepsilon}\in B$. Since by assumption $\{f\}_{n}>0$ and $f$ is extremal, from formula (8) for $\varepsilon<0$ we get
$\operatorname{Re}{(\{f\}_{n}\{h\}_{0}+\{f\}_{n-1}\{h\}_{1}+\ldots+\{f\}_{0}\{h\}_{n})}\leqslant 0,$
while for $\varepsilon>0$ formula (9) holds, i.e.,
$\operatorname{Re}{(\{f\}_{n}\{h\}_{0}+\{f\}_{n-1}\{h\}_{1}+\ldots+\{f\}_{0}\{h\}_{n})}\geqslant 0,$
from which we conclude that formula (10) is valid.
3. Let us prove (12). Fix $\zeta\in\bar{\Delta}$ and set
$g(z)\coloneq\frac{1+\zeta z}{1-\zeta z}=1+2\zeta z+2\zeta^{2}z^{2}+\ldots$
Clearly, $g\in C_{1}.$ Substituting the coefficients of $g$ into formula (9), we obtain
$\operatorname{Re}{(\{f\}_{n}+2\{f\}_{n-1}\zeta+\ldots+2\{f\}_{0}\zeta^{n})}\geqslant 0,\qquad|\zeta|\leqslant 1,$
which is equivalent to $\operatorname{Re}{H(z)}\geqslant 0$, $z\in\bar{\Delta}$. Since $H(0)>0$, by the open mapping principle $\operatorname{Re}H(z)>0$, $z\in\Delta$, which is equivalent to (12).
4. Let us prove formula (13). Fix $k$ to choose a specific $\varphi_{k}$ partially defining $h$ (see formula (5)) and consider the function
$g(z)\coloneq\frac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}=1+2e^{i\varphi_{k}}z+2e^{2i\varphi_{k}}z^{2}+\ldots$
Then $v=f\cdot e^{-\varepsilon g}\in B$ both for $\varepsilon\geqslant 0$ and for sufficiently small $\varepsilon<0$. That is, for $|\varepsilon|\leqslant\alpha_{k}$, where $\alpha_{k}$ partially defines $h$ (see formula (5)). From Lemma 1, the condition $\{f\}_{n}>0$, and the extremality of $f$, it follows that for $\varepsilon<0$
$\operatorname{Re}{(\{f\}_{n}+\{f\}_{n-1}2e^{i\varphi_{k}}+\ldots+\{f\}_{0}2e^{ni\varphi_{k}})}\leqslant 0,$
while for $\varepsilon\geqslant 0$
$\operatorname{Re}{(\{f\}_{n}+\{f\}_{n-1}2e^{i\varphi_{k}}+\ldots+\{f\}_{0}2e^{ni\varphi_{k}})}\geqslant 0,$
[…]
Theorem 3.
Let $n\in\mathbb{N}$, $f\coloneq e^{-h}$, where $h(z)\coloneq\sum\limits_{k=1}^{m}\alpha_{k}\dfrac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}$, $\alpha_{k}>0$, $k=1,\ldots,m$, $0\leqslant\varphi_{1}<\ldots<\varphi_{m}<2\pi$, $m\leqslant n$ and let $g(z)\coloneq\sum\limits_{k=1}^{m}\beta_{k}\dfrac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}$, $\beta_{k}>0$, $k=1,\ldots,m$, and let the polynomial $H$ be generated from the coefficients of $f$ by formula (11), with $H\in C$. The equality
$\operatorname{Re}{(\{f\}_{n}\{g\}_{0}+\{f\}_{n-1}\{g\}_{1}+\ldots+\{f\}_{0}\{g\}_{n})}=0$
holds if and only if all equalities
$\operatorname{Re}{H(e^{i\varphi_{k}})}=0,\qquad k=1,\ldots,m.$
hold.
Proof. We have
$g(z)=\sum\limits_{k=1}^{m}\beta_{k}\dfrac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}=\sum\limits_{k=1}^{m}\beta_{k}(1+2e^{i\varphi_{k}}z+\ldots+2e^{in\varphi_{k}}z^{n}+\ldots).$
Applying formula (6), we obtain:
$\{f\cdot g\}_{n}=\sum\limits_{k=1}^{m}\beta_{k}(\{f\}_{n}+\{f\}_{n-1}2e^{i\varphi_{k}}+\ldots+\{f\}_{0}2e^{in\varphi_{k}})=\sum\limits_{k=1}^{m}\beta_{k}H(e^{i\varphi_{k}}).$ (14)
Since $H\in C$ by assumption, i.e., $\operatorname{Re}{H(z)}>0$, $z\in\Delta$ and $\operatorname{Re}{H(z)}\geqslant 0$, $z\in\partial\Delta$, it follows from formula (14) that
$\operatorname{Re}{\{f\cdot g\}_{n}}=0\quad\Leftrightarrow\quad\sum\limits_{k=1}^{m}\beta_{k}\operatorname{Re}{H(e^{i\varphi_{k}})=0}\quad\Leftrightarrow\quad\operatorname{Re}{H(e^{i\varphi_{k}})}=0,\leavevmode\nobreak\ k=1,\ldots,m.$
[…]
From Theorem 3 immediately follows:
Corollary 2.
Let $n\in\mathbb{N}$, $f$ be a function extremal in the Krzyz problem, and let $g(z)\coloneq\sum\limits_{k=1}^{n}\beta_{k}\dfrac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}$, $\beta_{k}>0$, $k=1,\ldots,n$, $0\leqslant\varphi_{1}<\ldots<\varphi_{n}<2\pi$. Then, if the function $g$ satisfies equality (10), i.e.,
$\operatorname{Re}{(\{f\}_{n}\{g\}_{0}+\ldots+\{f\}_{0}\{g\}_{n})}=0,$
then
$f(z)=\exp\left(-\sum_{k=1}^{n}\alpha_{k}\frac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}\right).$
That is, $m=n$ and the numbers $\varphi_{k}$, $k=1,\ldots,n$, are known.
Proof. According to Corollary 1, there exist numbers $\alpha_{k}>0$, $k=1,\ldots,m$, and $0\leqslant\theta_{1}<\ldots<\theta_{m}<2\pi$, where $m\leqslant n$, such that $h(z)=\sum\limits_{k=1}^{m}\alpha_{k}\dfrac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}$ and $f=e^{-h}$.
[…]
3. Riesz-Fejer theorem
Let $n\in\mathbb{N}$. According to Theorem 2, the polynomial $H$, generated from the first $n+1$ coefficients of the function extremal in the Krzyz problem for the coefficient with index $n$, has a positive real part in the disk $\Delta$. Let us derive some properties of polynomial elements of class $C$, necessary for studying the extremal function.
A function of a real argument $\varphi$
$T(\varphi)\coloneq a_{0}+\sum\limits_{k=1}^{n}(a_{k}\cos{k\varphi}-b_{k}\sin{k\varphi}),\qquad a_{0},a_{k},b_{k}\in\mathbb{R},\quad k=1,\ldots,n,$
is called a trigonometric polynomial of degree $n$ if the condition $a_{n}^{2}+b_{n}^{2}>0$ is satisfied. Obviously, the following statement holds:
Lemma 2.
Let a polynomial $H(z)\coloneq h_{0}+2\sum\limits_{k=1}^{n}h_{k}z^{k}$ be given, where
$h_{0}\coloneq a_{0}\in\mathbb{R},\qquad h_{k}\coloneq\frac{a_{k}+ib_{k}}{2},\quad a_{k},b_{k}\in\mathbb{R},\quad k=1,\ldots,n,\quad h_{n}\neq 0.$
Then the restriction of $\operatorname{Re}H$ to the unit circle $\partial\Delta$ is a trigonometric polynomial
$T(\varphi)\coloneq\operatorname{Re}H(e^{i\varphi})=a_{0}+\sum\limits_{k=1}^{n}(a_{k}\cos{k\varphi}-b_{k}\sin{k\varphi}).$
Moreover, if $H\in C$, then $T(\varphi)\geqslant 0$ for all $\varphi\in\mathbb{R}$.
As noted in the article [21], in the early 19th century L. Fejer [22] first drew attention to the importance of the class of trigonometric polynomials taking non-negative values on the entire real line. His conjecture about the structure of such polynomials was later proved by F. Riesz [23] and is now known as the Riesz-Fejer theorem.
Theorem 4 (Riesz, Fejer).
If $T(\varphi)\coloneq\sum\limits_{k=0}^{n}(a_{k}\cos{k\varphi}-b_{k}\sin{k\varphi})\geqslant 0$, $a_{n}^{2}+b_{n}^{2}>0$, $\varphi,a_{k},b_{k}\in\mathbb{R}$, then there exists a polynomial $P(z)\coloneq\sum\limits_{k=0}^{n}p_{k}z^{k}$, $p_{n}\neq 0$ such that $T(\varphi)=|P(e^{i\varphi})|^{2}$. Moreover, $P$ can be chosen so that all its zeros lie outside the unit disk. In the latter case, $P$ is uniquely determined up to a unimodular multiplicative constant.
The proof of the Riesz-Fejer theorem is given, in particular, in the article [24] and in the monograph [25, p. 154]. From Theorem 4 it follows that if $m$ out of $n$ roots of the polynomial $P$ lie on the unit circle, then the trigonometric polynomial $T$ has $2m$ roots.
Corollary 3.
A trigonometric polynomial of degree $n$ has at most $2n$ roots on $[0,2\pi)$ counting multiplicities.
4. Polynomials with positive real part
Let us present some theorems about harmonic functions that we will need for proving the necessary condition of Theorem 8. All formulations are adapted for the case of the unit disk $\Delta$.
Theorem 5.
For any harmonic function $g$ in $\Delta$, one can construct a holomorphic function $h$ in $\Delta$ for which $g$ is the real part. Moreover, $h$ is unique up to an additive purely imaginary constant.
Theorem 6 (Uniqueness theorem).
If two harmonic functions $g_{1}$ and $g_{2}$ in $\Delta$ coincide on a set $A\subset\Delta$ that has at least one interior point, then $g_{1}\equiv g_{2}$ in $\Delta$.
Note that the uniqueness theorem for harmonic functions imposes stricter conditions than the analogous theorem for holomorphic functions, where it is only required that the set $A$ contains a limit point. For example, the function $u(z)\coloneq x$, where $x\coloneq\operatorname{Re}z$, equals zero on the imaginary axis, but $u(z)\not\equiv 0$.
Theorem 7 (Solution to the Dirichlet problem for a disk).
If a function $g$ is continuous on $\partial\Delta$, then there exists a unique harmonic function $h$ in $\Delta$, coinciding with $g$ on $\partial\Delta$.
Theorems 56 and 7 are taken from the book [26] (pp. 239, 240 and 246 respectively).
Theorem 8.
A polynomial $H(z)\coloneq h_{0}+2\sum\limits_{k=1}^{n}h_{k}z^{k}$, $h_{0}>0$, has positive real part in $\Delta$ if and only if there exist numbers $p_{k}\in\mathbb{C}$, not all equal to zero, such that
$h_{k}=\sum\limits_{j=0}^{n-k}p_{j+k}\bar{p}_{j},\qquad k=0,\ldots,n.$ (15)
Proof. Fix $n\in\mathbb{N}$ and an arbitrary polynomial $P(z)\coloneq\sum\limits_{k=0}^{n}p_{k}z^{k}$. Consider the polynomial $Q(z)\coloneq q_{0}+2\sum\limits_{k=1}^{n}q_{k}z^{k}$, where $q_{k}\coloneq\sum\limits_{j=0}^{n-k}p_{j+k}\bar{p}_{j}$, $k=0,\ldots,n$. For all $\varphi\in\mathbb{R}$ we have:
$0\leqslant$ $\operatorname{Re}H(e^{i\varphi})=T(\varphi)=|P(e^{i\varphi})|^{2}=P(e^{i\varphi})\overline{P(e^{i\varphi})}=\sum\limits_{k=0}^{n}p_{k}e^{ik\varphi}\sum\limits_{k=0}^{n}\bar{p}_{k}e^{-ik\varphi}=$ (16)
$=$ $|p_{0}|^{2}+\ldots+|p_{n}|^{2}+$
$+(p_{0}\bar{p}_{1}+p_{1}\bar{p}_{2}+\ldots+p_{n-1}\bar{p}_{n})e^{-i\varphi}+(p_{1}\bar{p}_{0}+p_{2}\bar{p}_{1}+\ldots+p_{n}\bar{p}_{n-1})e^{i\varphi}+$
$+(p_{0}\bar{p}_{2}+p_{1}\bar{p}_{3}+\ldots+p_{n-2}\bar{p}_{n})e^{-i2\varphi}+(p_{2}\bar{p}_{0}+p_{3}\bar{p}_{1}+\ldots+p_{n}\bar{p}_{n-2})e^{i2\varphi}+$
$\ldots$
$+p_{0}\bar{p}_{n}e^{-in\varphi}+p_{n}\bar{p}_{0}e^{in\varphi}=$
$=$ $|p_{0}|^{2}+\ldots+|p_{n}|^{2}+$
$+2\operatorname{Re}(\{p_{1}\bar{p}_{0}+p_{2}\bar{p}_{1}+\ldots+p_{n}\bar{p}_{n-1}\}e^{i\varphi})+$
$+2\operatorname{Re}(\{p_{2}\bar{p}_{0}+p_{3}\bar{p}_{1}+\ldots+p_{n}\bar{p}_{n-2}\}e^{i2\varphi})+$
$\ldots$
$+2\operatorname{Re}(p_{n}\bar{p}_{0}e^{in\varphi})=$
$=$ $\operatorname{Re}{Q(e^{i\varphi})}.$
Necessity. Let $H\in C$. We show that $H$ has coefficients (15). By Lemma 2, the expression $T(\varphi)\coloneq\operatorname{Re}H(e^{i\varphi})$ is a trigonometric polynomial of degree $n$, and $T(\varphi)\geqslant 0$ for all $\varphi\in\mathbb{R}$, since $H\in C$. According to Theorem 4, there exists a polynomial $P$ such that $T(\varphi)=|P(e^{i\varphi})|^{2}$. Thus, from (16) it follows that $\operatorname{Re}H(e^{i\varphi})=\operatorname{Re}Q(e^{i\varphi})$.
Let us perform the inverse substitution $e^{i\varphi}=z$. We have $\operatorname{Re}H(z)\equiv\operatorname{Re}Q(z)$, $z\in\partial\Delta$. According to Theorem 6, this does not mean that the functions $\operatorname{Re}H(z)$ and $\operatorname{Re}Q(z)$ coincide everywhere in $\Delta$. However, by Theorem 7 these functions can be uniquely continued to harmonic functions in $\Delta$. Let us reconstruct the analytic functions $Q$ and $H$ from their real parts which coincide in $\Delta$ (this can be done according to Theorem 5). Taking into account the normalization $Q(0)=H(0)>0$, we conclude that $H\equiv Q$ in $\Delta$, since a holomorphic function in $\Delta$ is reconstructed from its real part uniquely up to a purely imaginary additive constant by Theorem 5. As required.
[…]
5. Uniqueness conditions for polynomials from the Caratheodory class
In this section we assume that $n\in\mathbb{N}$, $h_{0}>0$ and
$H(z)\coloneq h_{0}+2\sum\limits_{k=1}^{n}h_{k}z^{k}.$ (17)
Lemma 3.
There exists a unique polynomial $H$ of the form (17), up to rotations in the $z$-plane, satisfying:
  • $H\in C$;
  • $h_{0}=2|h_{n}|>0$.
Moreover, $H(z)=h_{0}(1+\eta z^{n})$, where $|\eta|=1$, and $\arg\eta=\arg\{h\}_{n}$.
Proof. By Theorem 8, there exist numbers $p_{k}\in\mathbb{C}$, $k=0,\ldots,n$, satisfying:
$h_{0}=|p_{0}|^{2}+\ldots+|p_{n}|^{2},\qquad h_{n}=p_{n}\bar{p}_{0}.$
From $h_{0}=2|h_{n}|$ we obtain:
$|p_{0}|^{2}+\ldots+|p_{n}|^{2}=2|p_{n}p_{0}|.$
This equality is equivalent to:
$(|p_{0}|^{2}-2|p_{n}p_{0}|+|p_{n}|^{2})+|p_{1}|^{2}+\ldots+|p_{n-1}|^{2}=0,$
which can be rewritten as:
$(|p_{0}|-|p_{n}|)^{2}+|p_{1}|^{2}+\ldots+|p_{n-1}|^{2}=0.$
This immediately implies $|p_{0}|=|p_{n}|$ and $p_{1}=\ldots=p_{n-1}=0$. Thus,
$H(z)=2|p_{0}|^{2}+2p_{n}\bar{p}_{0}z^{n}=2|p_{0}|^{2}(1+\eta z^{n}),\qquad|\eta|=1,\quad\arg\eta=\arg\{h\}_{n}.$
[…]
Note that the roots $z_{k}$ of $H(z)=h_{0}(1+z^{n})$ are exactly all $n$-th roots of $-1$, i.e. $z_{k}=e^{i\frac{\pi+2\pi k}{n}}$, $k=0,\ldots,n-1$.
Statement 1.
There exists a unique polynomial $H$ of the form (17) satisfying:
  • $H\in C$;
  • $h_{n}>0$;
  • All roots $z_{1},\ldots,z_{n}$ of $H$ lie on the unit circle $\partial\Delta$.
Moreover, $H(z)=h_{0}(1+z^{n})$.
[…]
Statement 2.
There exists a unique polynomial $H$ of the form (17) satisfying:
  • $H\in C$;
  • $h_{n}>0$;
  • All $n$-th roots of $-1$ and only they are roots of $\operatorname{Re}H$.
Moreover, $H(z)=h_{0}(1+z^{n})$.
[…]
6. Theorem on uniqueness of extremal function
Under the constraints we imposed on the extremal function, the question of its uniqueness can be raised. The following result relates to the main result of [16, p. 735].
Theorem 9.
Let $n\in\mathbb{N}$, and let $f$ be an extremal function in the Krzyz problem satisfying the condition $\{f\}_{n}=2\{f\}_{0}>0$. Then the function $f$ exists, is unique, and $f=F(z^{n},1)$, where the function $F$ is given by formula (1).
Proof. According to Theorem 2, the polynomial $H$, generated from the coefficients of $f$ by formula (11), belongs to class $C$ and satisfies all conditions of Lemma 3. Consequently,
$H(z)=\{f\}_{n}+2\{f\}_{0}z^{n},$
from which it follows that
$\{f\}_{1}=\ldots=\{f\}_{n-1}=0.$
Let $\{f\}_{0}=e^{-t}$, where $t>0$. Then, since by condition $\{f\}_{n}=2\{f\}_{0}$, we have:
$f(z)=e^{-t}+2e^{-t}z^{n}+o(z^{n}).$
Let us verify that the function $h\coloneq-\ln f=t-2z^{n}+o(z^{n})$ belongs to class $C$. (Recall that the minors $M_{k}$ are defined by formula (4).) Applying the Caratheodory-Toeplitz criterion, we obtain for $t>0$:
  • Minors $M_{k}=2^{k+1}t^{k+1}>0$ for $k=1,\ldots,n-1$;
  • $M_{n}=2^{n+1}t^{n-1}(t^{2}-1)\geqslant 0$ and $M_{n}=0$ if and only if $t=1$.
Thus, $\{f\}_{0}=e^{-1}$ and $f(z)=Q_{n}(z)+o(z^{n})$, where $Q_{n}(z)\coloneq e^{-1}+2e^{-1}z^{n}$.
[…]
7. A proof of the Krzyz conjecture
Theorem 10.
Let $n\in\mathbb{N}$ and $f\in B$. Then the following inequality holds:
|$|\{f\}_{n}|\leqslant\frac{2}{e};$
equality in this inequality is achieved if and only if
$f(z)=e^{i\theta}F(e^{i\varphi}z^{n},1),\quad\varphi,\theta\in\mathbb{R},$
where the function $F$ is defined by formula (1).
Proof. Fix an arbitrary index $n$ and an arbitrary extremal function $f\in B$. According to Corollary 1, there exist numbers:
  • $\alpha_{k}>0$, $k=1,\ldots,m$, $\sum_{k=1}^{m}\limits\alpha_{k}=t$, where $m\leqslant n$, and $t\coloneq-\ln\{f\}_{0}$;
  • $0\leqslant\varphi_{1}<\ldots<\varphi_{m}<2\pi$;
such that the function $f$ can be represented as:
$f(z)=\exp\left(-\sum_{k=1}^{n}\alpha_{k}\frac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}\right).$
Direct verification shows that the function $F(z^{n},1)=e^{-1}+2e^{-1}z^{n}+o(z^{n})$ satisfies [^2] Error here: not $F$ and $-\ln F$ should satisfy (10), but $f$ and $-\ln F$. relation (10), therefore, from Corollary 2 it follows that: $m=n$, the points $e^{i\varphi_{k}}$, $k=1,\ldots,n$, are all $n$-th roots of $-1$, since $F(z^{n},1)=\exp\left(-\dfrac{1-z^{n}}{1+z^{n}}\right)=\exp\left(-\dfrac{1}{n}\sum\limits_{k=1}^{n}\dfrac{1+e^{i\varphi_{k}}z}{1-e^{i\varphi_{k}}z}\right).$
Consider the polynomial $H$ generated from the coefficients of $f$ by formula (11). From Theorem 2 it follows that:
  • $H\in C$,
  • $\operatorname{Re}H(e^{i\varphi_{k}})=0$, $k=1,\ldots,n$.
By Statement 2, there exists a unique polynomial $H\in C$ of degree $n$, whose real part has as its roots all possible $n$-th roots of $-1$, where:
$H(z)=\{f\}_{n}+2\{f\}_{0}z^{n}=h_{0}(1+z^{n}),$
with $h_{0}$ being a positive constant factor. Thus,
$\{f\}_{n}=2\{f\}_{0}.$
[…]
Conclusion
Problems of geometric theory of functions of a complex variable, in one way or another related to the Krzyz conjecture, are considered in works [27, 28, 7, 29, 30, 31, 15, 32, 33]. Some of these problems are equivalent to the Krzyz conjecture and, therefore, can also be considered as solved. Some generalizations of the Krzyz conjecture are described in [4, p. 187].